OpenStudy (anonymous):

How do I use Distributive Prop. or Mental Math to do: 5(42) basicly 5x42=

OpenStudy (anonymous):

yes 42*5 = 210

OpenStudy (anonymous):

yes its correct...

OpenStudy (anonymous):

Ok so just do 5x42?

OpenStudy (asnaseer):

what you should do is as follows: 5(42) = 5(40 + 2) = 5*40 + 5*2 = 200 + 10 = 210

OpenStudy (anonymous):

yes ignore the parentheses its still the same as multiplication

OpenStudy (anonymous):

Ok thank you!

OpenStudy (anonymous):

he provided another a good point too^

OpenStudy (asnaseer):

the idea is to split the multiplication up into simpler components

OpenStudy (anonymous):

Oh ok thank you that helps alot!! :)

OpenStudy (asnaseer):

yw :)